OpenStudy (anonymous):

A doctor’s office schedules 15-minute appointments and half-hour appointments for weekdays. The doctor limits these appointments to, at most, 30 hours per week. Write an inequality to represent the number of 15-minute appointments, x, and the number of half-hour appointments, y, the doctor may have in a week. Please be patient for my replies, I type slow.

OpenStudy (anonymous):

30 hours = 30 * 60 = 1800 mins 15*x + 30 * y <= 1800

OpenStudy (anonymous):

Wow, that is a lot simpler than I thought...thanks linkinage!